OpenStudy (anonymous):

A line passes through the point (-4,7) and has a slope of -3. What is the equation for this line?

OpenStudy (mathlegend):

@zippers you need to give explanations.

OpenStudy (mathlegend):

Not just the answer.

OpenStudy (mathlegend):

@cici123 to show some work... I would do this.

OpenStudy (mathlegend):

y-y1=m(x-x1)

OpenStudy (mathlegend):

They gave a slope, y1 and a x1

OpenStudy (mathlegend):

y-7=-3(x+4) y-7=-3x-12 +7 +7 ------------ y=-3x-5

OpenStudy (mathlegend):

@cici123 do you understand?

OpenStudy (anonymous):

yes thank you!

OpenStudy (mathlegend):

:)